m<rxq =(x+15)° m<rxs=(5x-7)° m<qxs =(3x+5)° find the x value and m<rxs


1
Expert's answer
2021-01-20T17:53:30-0500

Solution:

"m<RXQ+m<QXS=m<RXS"

"x+15+3x+5=5x-7"

"4x+20=5x-7"

"4x-5x=-20-7"

"-1x=-27"

"x=27"

"m<RXS=5x-7=5*27-7=128\degree"

Answer:

"x=27" , "m<RXS=128\degree"